uzsbgck:

261 grams to kg ( Use scientific notation and units to answer)

AZ:

 So we have 261.0 How many times are we moving over the decimal point?

uzsbgck:

2?

AZ:

Yes, so each time we move the decimal point back we're pulling out a factor of 10 So that means we were going from 261 = 26.1 * 10 = 2.61 * 10 * 10

AZ:

Scientific notation is writing a number as \(a \times 10^n\) and a has to be a number between 0 and 9

AZ:

How can you write 10*10 as an exponent?

uzsbgck:

wait so what would the unit be after that

AZ:

oh you're right, that completely slipped my mind I was just trying to write it in scientific notation

AZ:

Do you know how many grams make up 1 kilogram?

uzsbgck:

1000

AZ:

Yes! So to convert from grams to kg, we have to divide by 1000 What is 261/1000 = ?

uzsbgck:

0.261

AZ:

so 261 grams = 0.261 kg Now w need to write 0.261 kg in scientific notation

AZ:

Remember what I said earlier \(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ Scientific notation is writing a number as \(a \times 10^n\) and a has to be a number between 0 and 9 \(\color{#0cbb34}{\text{End of Quote}}\)

AZ:

We need to write 0.261 as 2.61 * 10^something Does that make sense so far?

uzsbgck:

yes

AZ:

 We have to move the decimal one over to the right to make it 2.61 So in case you don't know 10^(-1) = 1/10 = 0.1 10^0 = 1 10^1 = 10 So we have 0.261 * 10^0 and we're moving the decimal point to the left, so the exponent will decrease by 1 so 2.61 * 10^what what is 0-1? This image might help you see it better too

1 attachment
uzsbgck:

0 - 1 = -1

AZ:

so 0.261 * 10^0 is the same as 2.61 * 10^(-1)

AZ:

And that's your answer in kg written in scientific notation

uzsbgck:

Ok and do i write Kg after that for the unit since it was converted?

AZ:

Yes!

uzsbgck:

ok thank you so much

AZ:

It was my pleasure!